OpenStudy (anonymous):

Write the equation for the line which passes through the point (-1, 5) and is perpendicular to the graph of y = 2x + 6. Write the equation in slope-intercept form. Answer y = 1/2x + 11/2 y = -1/2x + 11/2 y = -1/2x + 9/2 y = -1/2x + 1/2

OpenStudy (oaktree):

Well, first, to find a line, we need the slope.

OpenStudy (oaktree):

If the line is perpendicular to a line of slope 2, then the slope of OUR line is the negative reciprocal, or -1/2.

OpenStudy (oaktree):

Good until now?

OpenStudy (oaktree):

I'll keep going, then!

OpenStudy (oaktree):

Next, we can use our coordinates and the slope to get our y-intercept. Plugging in, we get 5=-(-1/2)+b --> 5=1/2+b --> b=5-1/2=10/2-1/2=9/2.

OpenStudy (oaktree):

So the line is y=-1/2x+9/2.

OpenStudy (anonymous):

thanks lol